# Recent Advances in Discrete Speech Tokens: A Review

### v2

@@ V-D Voice Conversion Analysis
Despite the loss of acoustic information, a prominent advantage of semantic tokens over most acoustic tokens is their inherent timbre controllability. Some acoustic tokens also have this ability, such as those with a global encoder like TiCodec and disentangled acoustic tokens, also possess this ability To compare this ability across different tokens, we conduct voice conversion (VC) experiments using these tokens as the content from the source speech. We use the same source utterances in Section V-C, but assign a different target speaker for each source utterance as the prompt. Then, we perform VC experiments on the 500 VC pairs. In addition to WER and SECS, we also measure P.Corr as an objective metric for prosody preservation. This calculates the Pearson correlation coefficient between the pitch contours of the converted and source utterances. Note that this metric will be meaninglessly high if the VC similarity is low, i.e., when the source timbre is barely changed. As the source utterances are the same as Section V-C, the WER numbers are directly comparable to those in the reconstruction experiments.
@@
The results presented in Table III indicate that semantic tokens often achieve much higher VC similarity compared to acoustic tokens. However, due to the substantial loss of prosody information, semantic tokens tend to have lower P.Corr scores than acoustic tokens. Among the acoustic tokens capable of performing VC, explicit disentanglement methods, such as FACodec and LSCodec, outperform the implicit criterion employed in TiCodec. It is also noteworthy that different tokenization settings in wav2vec 2.0 lead to drastically different outcomes. Tokens generated from its inner quantizer preserve prosody well but also retain much speaker information, whereas clusters derived from its Transformer hidden embeddings exhibit the opposite characteristics.
@@
Supervised semantic tokens from $\mathcal{S}^{3}$ Tokenizer also exhibit good intelligibility and VC ability. Unlike HuBERT-style SSL models, this supervised tokenizer demonstrates better preservation of prosody both in reconstruction and VC settings. Given that prosody and intonation are a crucial factors for ASR, it is reasonable to assume that the tokenizer’s VQ module encodes some prosody information. In contrast, while HuBERT-style SSL models do contain rich prosody information in their continuous features (e.g., as evidenced by good emotion recognition results [117]), phonetic information is likely the primary component. Therefore, offline clustering is prone to discard these prosody characteristics.
VI Discrete Speech Token-based Application Paradigms
VI-A Spoken Language Understanding
VI-A1 Motivation
@@
VI-A2 Speech Translation
Among the various SLU tasks, discrete speech tokens are mostly adopted in speech translation, including speech-to-text translation (S2TT) and speech-to-speech translation (S2ST). Since semantic tokens correlate well with phonetics, they can serve as universal pseudo-labels for untranscribed languages, useful for S2TT in low-resource settings [187]. Direct S2ST using discrete tokens has garnered more attention on the generation side (Section VI-B). Early approaches primarily rely on extracting discrete tokens using VQ-VAEs, particularly for unwritten languages [188, 189]. Recent researches in this area include employing semantic tokens [190, 191, 192], acoustic tokens [193, 194, 195], two-pass architectures [196, 197], and non-autoregressive frameworks [198]. These efforts collectively contribute to advancing the performance and applicability of discrete token-based speech translation systems.
VI-A3 Unified Speech Understanding
Discrete tokens provide opportunity to construct unified and adaptable spoken language models with various SLU functionalities. Efforts include task identifies [7], prompt tuning [199, 200, 201, 192], shared audio and text vocabulary [202], and combining continuous and discrete representations [203]. These efforts highlight the potential of discrete tokens in enhancing the performance and versatility of universal SLU models.
VI-A4 Limitations
Despite the advantages and growing popularity in S2ST tasks, discrete tokens still underperform in many SLU tasks. Lots of SLU studies [204, 156, 205, 180, 181] only verify that discrete tokens can surpass traditional frequency-domain features in certain tasks such as ASR. Continuous SSL features continue to have superior performance [182]. The majority of current LLM-based SLU models rely predominantly on continuous inputs, such as Whisper features [206, 207, 208, 209, 210, 211]. Moreover, the performance of discrete tokens in speaker-related tasks is generally much inferior to that of continuous features [204, 144]. A significant limitation of discrete tokens for SLU is the inevitable information loss during the quantization process. Mitigating such loss with more VQ codebooks may hinder the accessibility of semantic information crucial for SLU as well. Therefore, the full potential of leveraging discrete tokens for SLU remains largely untapped and warrants further exploration.
VI-B Speech Generation
VI-B1 Motivation
@@
  • Easier training objectives. Discrete tokens replace the original spectrogram-based regression task with a classification task [129], which can be much easier. This also offers a better balance between acoustic models and vocoders, since texts are closer to discrete speech tokens than frequency-domain features.
@@
  • Better use of decoder-only language models. Decoder-only language models have shown remarkable success in natural language generation. After discretization, speech can also be autoregressively generated under the same paradigm. This offers huge potential in leveraging the in-context learning and scaling capabilities of language models to achieve zero-shot high-fidelity TTS [6].
@@
Other generative tasks, such as singing voice synthesis and speech editing, can similarly benefit from the advantages of discrete tokens observed in TTS. For voice conversion (VC), using discrete tokens as content representations can simplify the process to a token vocoder [171], when timbre information is effectively removed from the tokens. Tasks like speech to speech translation [190, 189], speech enhancement [212, 213] and target speaker extraction [214] can also be enhanced through language modeling on discrete tokens.
VI-B2 Autoregressive TTS
@@
Autoregressively predicting the next VQ index of discrete speech tokens is first proposed in VQTTS [129], which uses an LSTM conditioned on Transformer representations to generate vq-wav2vec [31] semantic tokens. A discrete token vocoder converts the tokens to waveforms with the assistance of handcrafted prosody features. VQTTS achieves state-of-the-art TTS quality at that time, and shows promising performance in speaker-adaptive TTS [215, 216, 217] and expressive TTS [218].
@@
Subsequently, decoder-only TTS models using neural audio codecs have made tremendous success in zero-shot TTS starting from VALL-E [6]. VALL-E contains an autoregressive (AR) model and non-autoregressive (NAR) model, both of which generate EnCodec [63] RVQ tokens. The AR model performs next-token prediction on the first RVQ layer conditioned on text. The NAR model predicts the $n+1$-th RVQ tokens given the text, all EnCodec tokens from the speaker reference, and the previous $n$ RVQ layers. VALL-E employs a concise design in which text and speaker references serve as “prompts” for a language model. It achieves remarkable zero-shot TTS performance when trained on 60k hours of speech. Later, methods have been proposed to improve generation robustness [219, 220, 221, 222, 223, 224], efficiency [225], style control [226, 227, 228], and to incorporate LLMs [229, 230].
@@
Besides using an NAR model to predict the rest RVQ layers, alternate modeling strategies have been proposed, such as hierarchical modeling [231] and token interleaving patterns [232, 233]. Semantic tokens are also introduced to cooperate with acoustic codecs [15, 130, 230, 234], which might decrease the modeling difficulty since they bridge the gap between texts and acoustics and usually require only a single token stream. Numerous industry-level large-scale TTS systems have been produced in this autoregressive TTS paradigm, such as XTTS [235], BASE-TTS [236], Seed-TTS [173], CosyVoice [41, 153], Fish-Speech [237], etc.
VI-B3 Non-Autoregressive TTS
@@
While autoregressive modeling is the current mainstream of TTS with discrete tokens, non-autoregressive models also exist. These models either treat the code-vectors as continuous features [238], or directly generate discrete tokens by masked prediction [136] or discrete diffusion models [111, 28]. These non-autoregressive methods are naturally more robust than autoregressive methods in inference, and also supports speech editing.
VI-B4 Unified Speech Generation
@@
The language modeling approach of discrete tokens allows a unified generation framework for multiple tasks. It suffices to use a task identifier to condition the unified language model. For example, [239, 240] extends VALL-E with more tasks like cross-lingual TTS, S2ST, speech editing, etc. UniAudio [231] supports 11 speech and audio generation tasks within a single hierarchical Transformer model. Prompt tuning upon a spoken language model has also been explored in [192] for efficient, transferable and versatile generation. These efforts demonstrate the potential of a large-scale foundation model for generation.
VI-B5 Limitations
@@
In contrast to discrete tokens, another emerging framework for speech generation is diffusion or flow matching-based models, including non-autoregressive models [241, 242, 243, 244] or autoregressive models [245, 246, 247, 248, 249]. They generate continuous features, and some even eliminate the need for forced alignments in non-autoregressive generation. Owing to the strong capability of diffusion and flow matching algorithms, they also have remarkable generation fidelity, diversity and controllability. They can have a higher upper bound for speech quality and intelligibility, as they inherently avoid quantization errors. In comparison, discrete token-based speech generation models usually fall short in generation robustness. Therefore, there is an ongoing debate between discrete and continuous representations for speech generation.
VI-C Text-Free Spoken Language Models
VI-C1 Motivation
@@
VI-C2 Existing Efforts
Ever since GSLM [4] and AudioLM [5] proposed the vision of TF-SLMs, building such models remains a significant challenge even till today. This difficulty primarily arises from the lower language modeling efficiency of speech token sequences compared to text, due to their lower semantic information density, longer sequence lengths, and the presence of paralinguistic information [250]. Current advancements in TF-SLMs mainly focus on two strategies: (1) reducing token frame rates, and (2) aligning speech with text.
The first approach aims to shorten speech sequences and enhance semantic density by lowering frame rates [4, 131, 159] to even $\approx$5Hz [45, 168]. While mitigating sequence length issues to different degrees, they still encounter scalability limitations [251] and compromise reconstruction quality. The second strategy involves aligning speech with text through methods like initializing pre-training with text LLMs [131], reinforcement learning using ASR and text LLM feedback [252], text-speech token interleaving [253], adopting novel architectures applied in text language modeling [254], etc [255, 256]. Meanwhile, full duplex modeling has been proposed [257] to allow users to interrupt and start new dialogues at will. However, despite many efforts, these models still struggle to generate semantically reliable long speech during inference due to the lack of explicit transcription guidance.
VI-C3 Limitations
@@
VI-D Text-Guided Spoken Language Models
VI-D1 Motivation
Since TF-SLM remains an open problem, the prevalent successful speech dialogue systems settle for an alternative choice that uses text as explicit guidance. Recent researches, especially following work like OpenAI’s GPT-4o [^10] https://openai.com/index/hello-gpt-4o/, have focused on SLMs that combine three key capabilities: strong understanding of speech semantics, high-quality speech output, and low latency [258, 259, 260, 10, 261, 155, 262, 263, 264, 265, 266, 154, 267, 268, 269, 270]. We refer to them as text-guided spoken language models (TG-SLMs). Unlike TF-SLMs, while TG-SLMs utilize a unified LLM for seamless processing of user’s speech input and system’s speech output, they internally decompose the end-to-end speech dialogue process into two well-established sub-procedures: SLU powered by LLMs, and real-time TTS. The two sub-procedures are connected via text as an intermediary to stabilize the semantic coherency of the final output. The LLM first generates a textual response to the audio input, then synthesizes the speech token sequence in a streaming fashion. In a TG-SLM, the SLU sub-procedure usually uses continuous speech features as input since they preserve more acoustic details for understanding, while the TTS sub-procedure typically uses discrete speech tokens as output to better fit LLM autoregressive generation.
VI-D2 Speech Generation in TG-SLMs
To reduce modeling complexity and better align with the autoregressive generation paradigm of LLMs, TG-SLMs favor single-layer discrete speech tokens as direct LLM outputs. Existing works make use of either the first layer of an RVQ codec [258], single-codebook codec [268], or single-codebook supervised semantic token [154, 269]. Specific designs are introduced corresponding to the tokens, such as chain-of-modality [258], token interleaving to lower latency [154], two-stage decoding process [268], etc. To better rebuild the speech information with the help of pretrained LLMs, several TG-SLMs use multi-layer speech tokens as LLM output, such as [10, 260, 261]. They often employ different techniques to generate the text tokens and multi-layer speech tokens in parallel reduce latency.
Mainstream TG-SLMs with discrete tokens as LLM outputs need an additional decoder to synthesize continuous speech signals, either using the codec decoder or a separately-trained vocoder. There are also efforts to streamingly synthesize speech signals directly based on the LLM hidden embedding [262, 270], eliminating the need for discrete tokens, additional decoders, or even explicit text tokens, hence further improving the real-time ability.
VI-D3 Limitations

### v3

@@ VI-D Voice Conversion Analysis
Despite the loss of acoustic information, a prominent advantage of semantic tokens over most acoustic tokens is their inherent timbre controllability. Some acoustic tokens also have this ability, such as those with a global encoder like TiCodec and disentangled acoustic tokens, also possess this ability To compare this ability across different tokens, we conduct voice conversion (VC) experiments using these tokens as the content from the source speech. We use the same source utterances in Section VI-C, but assign a different target speaker for each source utterance as the prompt.
Then, we perform VC experiments on the 500 VC pairs. In addition to WER, we also measure SECS (speaker embedding cosine similarity) [casanova2022yourtts] as the metric for speaker similarity, and P.Corr [guo2024vec2wav] as an objective metric for prosody preservation. SECS requires a speaker verification model [^6] https://github.com/resemble-ai/Resemblyzer to output speaker embeddings. P.Corr calculates the Pearson correlation coefficient between the pitch contours of the converted and source utterances. Note that P.Corr will be meaninglessly high if the VC similarity is low, i.e., when the source timbre is barely changed. As the source utterances are the same as Section VI-C, the WER numbers are directly comparable to those in the reconstruction experiments.
@@
The results presented in Table I indicate that semantic tokens often achieve much higher VC similarity compared to acoustic tokens. However, due to the substantial loss of prosody information, semantic tokens tend to have lower P.Corr scores than acoustic tokens. Among the acoustic tokens capable of performing VC, explicit disentanglement methods, such as FACodec and LSCodec, outperform the implicit criterion employed in TiCodec. It is also noteworthy that different tokenization settings in wav2vec 2.0 lead to drastically different outcomes. Tokens generated from its inner quantizer preserve prosody well but also retain much speaker information, whereas clusters derived from its Transformer hidden embeddings exhibit the opposite characteristics.
@@
Supervised semantic tokens from $\mathcal{S}^{3}$ Tokenizer also exhibit good intelligibility and VC ability. Unlike HuBERT-style SSL models, this supervised tokenizer demonstrates better preservation of prosody both in reconstruction and VC settings. Given that prosody and intonation are a crucial factors for ASR, it is reasonable to assume that the tokenizer’s VQ module encodes some prosody information. In contrast, while HuBERT-style SSL models do contain rich prosody information in their continuous features (e.g., as evidenced by good emotion recognition results [superb]), phonetic information is likely the primary component. Therefore, offline clustering is prone to discard these prosody characteristics.
VI-E Downstream Semantic Modeling Analysis
Besides reconstruction and voice conversion which mainly cover acoustic aspects, we also explore the semantic modeling abilities in the tokens. To this end, we follow the DASB [mousavi2024dasb] setup and considers two representative downstream semantic modeling tasks: ASR and intent classification (IC). In both tasks, we train a small LSTM-based probing network to process the discrete speech tokens and produce outputs. IC is a classification task to determine the spoken intents from speech directly, where we use the SLURP [bastianelli-etal-2020-slurp] dataset with 18 target classes and accuracy (ACC) as the metric. For ASR, we use LibriSpeech [librispeech] 960h to train the probing network and report WER on test-clean split.
The results can be found at the last two columns of Table I. General-purpose acoustic tokens show relatively worse semantic modeling abilities than others under a small probing model, although some acoustic tokens have excellent reconstruction quality. Semantic distillation can enhance such downstream performance, especially X-Codec which uses SSL features both as input and outputs. Meanwhile, semantic tokens generally achieve significantly better results in such tasks, with predictive SSL tokens being the best. For semantic tokens, downstream ASR performance appears to be negatively correlated with prosody preservation: tokens that better preserve prosody often yield worse ASR results. This trend is observed in vq-wav2vec, wav2vec 2.0, and even the supervised $\mathcal{S}^{3}$ Tokenizer, all of which employ an internal quantizer. In summary, these experiments reveal a trade-off between how strongly semantic information is emphasized in the tokens, and how comprehensively it is preserved.
VII Discrete Speech Token-based Application Paradigms
VII-A Spoken Language Understanding
VII-A1 Motivation
@@
VII-A2 Speech Translation
Among the various SLU tasks, discrete speech tokens are mostly adopted in speech translation, including speech-to-text translation (S2TT) and speech-to-speech translation (S2ST). Since semantic tokens correlate well with phonetics, they can serve as universal pseudo-labels for untranscribed languages, useful for S2TT in low-resource settings [zhang-etal-2023-dub]. Direct S2ST using discrete tokens has garnered more attention on the generation side (Section VII-B). Early approaches primarily rely on extracting discrete tokens using VQ-VAEs, particularly for unwritten languages [tjandra2019speech, zhang2021uwspeech]. Recent researches in this area include employing semantic tokens [lee2022direct, lee-etal-2022-textless, wu2023speechgen], acoustic tokens [peng2024mslm, wang-etal-2024-speech, gong2024seamlessexpressivelm], two-pass architectures [chen-etal-2023-speech, inaguma-etal-2023-unity], and non-autoregressive frameworks [huang2023transpeech]. These efforts collectively contribute to advancing the performance and applicability of discrete token-based speech translation systems.
VII-A3 Unified Speech Understanding
Discrete tokens provide opportunity to construct unified and adaptable spoken language models with various SLU functionalities. Efforts include task identifies [wang2024viola], prompt tuning [chang2022speechprompt, chang2023speechprompt, 10.1109/TASLP.2024.3436618, wu2023speechgen], shared audio and text vocabulary [rubenstein2023audiopalm], and combining continuous and discrete representations [chen2023lauragpt]. These efforts highlight the potential of discrete tokens in enhancing the performance and versatility of universal SLU models.
VII-A4 Limitations
Despite the advantages and growing popularity in S2ST tasks, discrete tokens still underperform in many SLU tasks. Lots of SLU studies [puvvada2024discrete, chang2024exploring, shon2024discreteslu, cui2024exploring_context, cui2024exploring] only verify that discrete tokens can surpass traditional frequency-domain features in certain tasks such as ASR. Continuous SSL features continue to have superior performance [wang2024comparative]. The majority of current LLM-based SLU models rely predominantly on continuous inputs, such as Whisper features [gong2023joint, chu2023qwen, tang2024salmonn, wavllm, ma2024embarrassingly, bai2024seed]. Moreover, the performance of discrete tokens in speaker-related tasks is generally much inferior to that of continuous features [puvvada2024discrete, mousavi2024dasb]. A significant limitation of discrete tokens for SLU is the inevitable information loss during the quantization process. Mitigating such loss with more VQ codebooks may hinder the accessibility of semantic information crucial for SLU as well. Therefore, the full potential of leveraging discrete tokens for SLU remains largely untapped and warrants further exploration.
VII-B Speech Generation
VII-B1 Motivation
@@
  • Easier training objectives. Discrete tokens replace the original spectrogram-based regression task with a classification task [VQTTS], which can be much easier. This also offers a better balance between acoustic models and vocoders, since texts are closer to discrete speech tokens than frequency-domain features.
@@
  • Better use of decoder-only language models. Decoder-only language models have shown remarkable success in natural language generation. After discretization, speech can also be autoregressively generated under the same paradigm. This offers huge potential in leveraging the in-context learning and scaling capabilities of language models to achieve zero-shot high-fidelity TTS [valle].
@@
Other generative tasks, such as singing voice synthesis and speech editing, can similarly benefit from the advantages of discrete tokens observed in TTS. For voice conversion (VC), using discrete tokens as content representations can simplify the process to a token vocoder [guo2024vec2wav], when timbre information is effectively removed from the tokens. Tasks like speech to speech translation [lee2022direct, zhang2021uwspeech], speech enhancement [wang2024selm, liu2024joint] and target speaker extraction [tang2024tselm] can also be enhanced through language modeling on discrete tokens.
VII-B2 Autoregressive TTS
@@
Autoregressively predicting the next VQ index of discrete speech tokens is first proposed in VQTTS [VQTTS], which uses an LSTM conditioned on Transformer representations to generate vq-wav2vec [vq-wav2vec] semantic tokens. A discrete token vocoder converts the tokens to waveforms with the assistance of handcrafted prosody features. VQTTS achieves state-of-the-art TTS quality at that time, and shows promising performance in speaker-adaptive TTS [TNVQTTS, DSETTS, limmits23] and expressive TTS [liu2024storytts].
@@
Subsequently, decoder-only TTS models using neural audio codecs have made tremendous success in zero-shot TTS starting from VALL-E [valle]. VALL-E contains an autoregressive (AR) model and non-autoregressive (NAR) model, both of which generate EnCodec [encodec] RVQ tokens. The AR model performs next-token prediction on the first RVQ layer conditioned on text. The NAR model predicts the $n+1$-th RVQ tokens given the text, all EnCodec tokens from the speaker reference, and the previous $n$ RVQ layers. VALL-E employs a concise design in which text and speaker references serve as “prompts” for a language model. It achieves remarkable zero-shot TTS performance when trained on 60k hours of speech. Later, methods have been proposed to improve generation robustness [song2024ella, xin2024rall, han2024vall, du2024vall, wang2024attention, chen2024vall], efficiency [song24b_interspeech], style control [kim2023sc, lyth2024natural, ji2024textrolspeech], and to incorporate LLMs [hao2023boosting, shen2024get].
@@
Besides using an NAR model to predict the rest RVQ layers, alternate modeling strategies have been proposed, such as hierarchical modeling [yang2024uniaudio] and token interleaving patterns [musicgen, voicecraft]. Semantic tokens are also introduced to cooperate with acoustic codecs [kharitonov2023speak, vectokspeech, shen2024get, yang2024interleaved], which might decrease the modeling difficulty since they bridge the gap between texts and acoustics and usually require only a single token stream. Numerous industry-level large-scale TTS systems have been produced in this autoregressive TTS paradigm, such as XTTS [casanova2024xtts], BASE-TTS [lajszczak2024base], Seed-TTS [seedtts], CosyVoice [du2024cosyvoice, cosyvoice2], Fish-Speech [liao2024fish], etc.
VII-B3 Non-Autoregressive TTS
@@
While autoregressive modeling is the current mainstream of TTS with discrete tokens, non-autoregressive models also exist. These models either treat the code-vectors as continuous features [shen2024naturalspeech2], or directly generate discrete tokens by masked prediction [wang2024maskgct] or discrete diffusion models [du2024unicats, facodec]. These non-autoregressive methods are naturally more robust than autoregressive methods in inference, and also supports speech editing.
VII-B4 Unified Speech Generation
@@
The language modeling approach of discrete tokens allows a unified generation framework for multiple tasks. It suffices to use a task identifier to condition the unified language model. For example, [vallex, wang2024speechx] extends VALL-E with more tasks like cross-lingual TTS, S2ST, speech editing, etc. UniAudio [yang2024uniaudio] supports 11 speech and audio generation tasks within a single hierarchical Transformer model. Prompt tuning upon a spoken language model has also been explored in [wu2023speechgen] for efficient, transferable and versatile generation. These efforts demonstrate the potential of a large-scale foundation model for generation.
VII-B5 Limitations
@@
In contrast to discrete tokens, another emerging framework for speech generation is diffusion or flow matching-based models, including non-autoregressive models [lee2023hierspeech++, le2024voicebox, liu2024e1, chen2024f5] or autoregressive models [meng2024autoregressive, liu2024autoregressive, zhu2024autoregressive, turetzky2024continuous, jia2025ditar]. They generate continuous features, and some even eliminate the need for forced alignments in non-autoregressive generation. Owing to the strong capability of diffusion and flow matching algorithms, they also have remarkable generation fidelity, diversity and controllability. They can have a higher upper bound for speech quality and intelligibility, as they inherently avoid quantization errors. In comparison, discrete token-based speech generation models usually fall short in generation robustness. Therefore, there is an ongoing debate between discrete and continuous representations for speech generation.
VII-C Text-Free Spoken Language Models
VII-C1 Motivation
@@
VII-C2 Existing Efforts
Ever since GSLM [lakhotia2021generative] and AudioLM [borsos2023audiolm] proposed the vision of TF-SLMs, building such models remains a significant challenge even till today. This difficulty primarily arises from the lower language modeling efficiency of speech token sequences compared to text, due to their lower semantic information density, longer sequence lengths, and the presence of paralinguistic information [wang2024whyspeech]. Current advancements in TF-SLMs mainly focus on two strategies: (1) reducing token frame rates, and (2) aligning speech with text.
The first approach aims to shorten speech sequences and enhance semantic density by lowering frame rates [lakhotia2021generative, hassid2024textually, shen2024acoustic] to even $\approx$5Hz [baade2024syllablelm, cho2024sylber]. While mitigating sequence length issues to different degrees, they still encounter scalability limitations [cuervo-marxer-2024-scaling] and compromise reconstruction quality. The second strategy involves aligning speech with text through methods like initializing pre-training with text LLMs [hassid2024textually], reinforcement learning using ASR and text LLM feedback [lin2024align], text-speech token interleaving [nguyen2024spirit], adopting novel architectures applied in text language modeling [park2024longform], etc [veluri2024beyond, zhang2024intrinsicvoice]. Meanwhile, full duplex modeling has been proposed [ma2024language] to allow users to interrupt and start new dialogues at will. However, despite many efforts, these models still struggle to generate semantically reliable long speech during inference due to the lack of explicit transcription guidance.
VII-C3 Limitations
@@
VII-D Text-Guided Spoken Language Models
VII-D1 Motivation
Since TF-SLM remains an open problem, the prevalent successful speech dialogue systems settle for an alternative choice that uses text as explicit guidance. Recent researches, especially following work like OpenAI’s GPT-4o [^7] https://openai.com/index/hello-gpt-4o/, have focused on SLMs that combine three key capabilities: strong understanding of speech semantics, high-quality speech output, and low latency [zhang2024speechgptgen, fu2024vita, Xie2024MiniOmniLM, kyutai2024moshi, xie2024mini2, fang2024llamaomni, yu2024salmonnomni, zhong2024lyra, chen2024slamomni, chen2024emova, wang2024freezeomni, zeng2024glm, zhang2024internlm, fu2025vita, luo2025openomni, chen2025minmo]. We refer to them as text-guided spoken language models (TG-SLMs). Unlike TF-SLMs, while TG-SLMs utilize a unified LLM for seamless processing of user’s speech input and system’s speech output, they internally decompose the end-to-end speech dialogue process into two well-established sub-procedures: SLU powered by LLMs, and real-time TTS. The two sub-procedures are connected via text as an intermediary to stabilize the semantic coherency of the final output. The LLM first generates a textual response to the audio input, then synthesizes the speech token sequence in a streaming fashion. In a TG-SLM, the SLU sub-procedure usually uses continuous speech features as input since they preserve more acoustic details for understanding, while the TTS sub-procedure typically uses discrete speech tokens as output to better fit LLM autoregressive generation.
VII-D2 Speech Generation in TG-SLMs
To reduce modeling complexity and better align with the autoregressive generation paradigm of LLMs, TG-SLMs favor single-layer discrete speech tokens as direct LLM outputs. Existing works make use of either the first layer of an RVQ codec [zhang2024speechgptgen], single-codebook codec [fu2025vita], or single-codebook supervised semantic token [zeng2024glm, luo2025openomni]. Specific designs are introduced corresponding to the tokens, such as chain-of-modality [zhang2024speechgptgen], token interleaving to lower latency [zeng2024glm], two-stage decoding process [fu2025vita], etc. To better rebuild the speech information with the help of pretrained LLMs, several TG-SLMs use multi-layer speech tokens as LLM output, such as [kyutai2024moshi, Xie2024MiniOmniLM, xie2024mini2]. They often employ different techniques to generate the text tokens and multi-layer speech tokens in parallel reduce latency.
Mainstream TG-SLMs with discrete tokens as LLM outputs need an additional decoder to synthesize continuous speech signals, either using the codec decoder or a separately-trained vocoder. There are also efforts to streamingly synthesize speech signals directly based on the LLM hidden embedding [yu2024salmonnomni, chen2025minmo], eliminating the need for discrete tokens, additional decoders, or even explicit text tokens, hence further improving the real-time ability.
VII-D3 Limitations